OpenStudy (anonymous):

I am trying to learn graphing for me to get smarter and smarter and this was my stop point where I just can't get it. If someone could explain to me how to do this. I will give you a medal,fan, and fan testimonial.

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

I'd like to help, but you'll need to postpone your question / problem first.

OpenStudy (anonymous):

well I just can't understand like graphing things i.e (rise over run) I get it mixed up where I do (run over rise). Also I get mixed up with my quadrants if someone could explain that to me it would be helpful and maybe a way to memorize it.

OpenStudy (anonymous):

also with Y and X intercept.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@♪Chibiterasu |

imqwerty (imqwerty):

alright lets take an example and then graph it :)

imqwerty (imqwerty):

okay lets start with this - \(x+y=1\)

imqwerty (imqwerty):

what do you think the graph of this equation will look like? :) a line a circle a parabola something else

OpenStudy (anonymous):

something else?

OpenStudy (anonymous):

cause a parabola is when x = 0 and y = 0 right?

OpenStudy (anonymous):

a line wouldn't really be something that equals one. When you add y + x

OpenStudy (anonymous):

and a circle I really don't think it would be

OpenStudy (anonymous):

am I correct?

OpenStudy (anonymous):

hello? Is anyone there?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

is anyone there?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@undeadknight26

OpenStudy (agent0smith):

You're asking to be taught a pretty broad concept. The fact you didn't recognize x+y=1 as a linear equation means you should probably start a bit more basic. Maybe something like Khan academy: Maybe this https://www.khanacademy.org/math/cc-eighth-grade-math/cc-8th-linear-equations-functions/8th-solutions-to-two-var-linear-equations/v/descartes-and-cartesian-coordinates Or slope https://www.khanacademy.org/math/cc-eighth-grade-math/cc-8th-linear-equations-functions/8th-slope/v/slope-intuition-example

OpenStudy (anonymous):

yea I learned a little bit more from watching it

OpenStudy (anonymous):

maybe you can test me on something

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (anonymous):

@ganeshie8 @DanJS @tkhunny @TheSmartOne

OpenStudy (anonymous):

@agent0smith reveal yourself

OpenStudy (anonymous):

You were here for like 30 minutes and never was typing

OpenStudy (agent0smith):

It was probably a glitch in OS. But I already posted what you should do. OpenStudy is better used as a place to enhance things you know, not learn an entire topic (unless it's something small/short).

OpenStudy (mathmale):

Lima, great that you're eager to learn and want to start with graphing. You did well mentioning two subtopics of graphing right up front (rise over run = slope of a straight line, and quadrants) on which you wanted to focus. Best to spend sufficient time on those topics so that you gain at least a basic understanding of them BEFORE moving on to other topics. If you keep on adding topics you want to learn, other users stop answering you because few have time to explain entire topics at once. So, back to the drawing board, so to speak. The slope of a straight line is defined as "rise over run," which is often written as m = slope = rise / run, or m= (change in y) over (change in x). A horizontal line has a slope of zero, as the graph neither increases nor decreases as you increase x. A slant line with a slope of m = 1 / 1 increases by exactly the same amount as x increases: